Which of the following acid can completely ionise in water?
HF or
H2S or
H2CO3 or
CH3COOH.
1
Expert's answer
2018-09-17T08:59:54-0400
H2S and H2CO3 are very weak acids and they are cannot be ionizes in water. CH3COOH are organic acid, which is are more weaker (Ka = 1.76*10^-5), than HF (Ka = 7.2*10^-4). Answer is HF.